A wheel is divided into 20 sections of 18 degree arc. Each section is number 0 to 5. Let X be a random variable of the section number that a wheel spin lands on.

A player pays $2 to spin and wins three times the number of dollars as the number the land on. For example, if they land on 2 they win 6 dollars. Let Y be a random variable that represents the value of their winnings. Y=3X-2

a. What is the expected value of Y?

b. Find VAR(5Y)
Answer #1

here as X follows discrete uniform distribution with parameter a =1 and b=20

hence E(X)=(a+b)/2=(1+20)/2=10.5

and Var(X)=((b-a+1)2-1)/12 =399/12

a)expected value of Y =E(3X-2)=3*E(X)-2=3*10.5-2=29.5

b)Var(5Y)=52Var(Y)=25*Var(3X-2)=25*32*Var(X)=25*9*399/12=7481.25